MICROWAVE SPECTRA AND STRUCTURES OF H_{2}S-CuCl AND H_{2}S-AgCl.

\underline{N. R. WALKER}, D. WHEATLEY AND A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

An FT-MW spectrometer coupled to a laser ablation source has been used to obtain the pure rotational spectra of H_{2}S-CuCl and H_{2}S-AgCl. Both molecules are generated via laser ablation (532 nm) of a metal rod in the presence of CCl_{4} and argon and are stabilized by supersonic expansion. Rotational constants (B_{0}, D_{J} ) have been measured for four isotopomers of each molecule with substitutions at the metal and chlorine atoms in each case. The spectrum of each molecule is consistent with a linear arrangement of sulphur, metal and chlorine atoms. Nuclear quadrupole coupling constants have been measured for the chlorine atom in each molecule and also for copper in H_{2}S-CuCl. Nuclear spinrotation constants have been determined for the copper atom in H_{2}S-CuCl. The observed trends are consistent with results previously reported for Ar-MX, OC-MX (where M is a Group 11 metal atom and X is a halogen) and N_{2}-CuF. Further experiments are being performed to obtain and assign the spectra of D_{2}S-CuCl and D_{2}S-AgCl. These data will allow determination of the angle defined between the axis of the D_{2}S molecule and that of the metal chloride unit.